### v1

marginparsep has been altered.
topmargin has been altered.
marginparpush has been altered.
The page layout violates the ICML style.Please do not change the page layout, or include packages like geometry, savetrees, or fullpage, which change it for you. We’re not able to reliably undo arbitrary changes to the style. Please remove the offending package(s), or layout-changing commands and try again.
The Flexibility Trap: Why Arbitrary Order Limits Reasoning Potential in Diffusion Language Models
Zanlin Ni $^{\,1\,}$, Shenzhi Wang 1, Yang Yue 1, Tianyu Yu 2, Weilin Zhao 2, Yeguo Hua 3, Tianyi Chen 3, Jun Song 4, Cheng Yu 4, Bo Zheng 4, Gao Huang $^{\,1\,\textrm{\Letter}}$
$^{1\,}$ LeapLab, Tsinghua University  $^{2\,}$ NLPLab, Tsinghua University
$^{3\,}$ Tsinghua University  $^{4\,}$ Alibaba Group
nzl22@mails.tsinghua.edu.cn, gaohuang@tsinghua.edu.cn
Diffusion Large Language Models (dLLMs) break the rigid left-to-right constraint of traditional LLMs, enabling token generation in arbitrary orders. Intuitively, this flexibility implies a solution space that strictly supersets the fixed autoregressive trajectory, theoretically unlocking superior reasoning potential for general tasks like mathematics and coding. Consequently, numerous works have leveraged reinforcement learning (RL) [^1] In this paper, unless otherwise specified, we use the term RL to refer specifically to Reinforcement Learning with Verifiable Rewards (RLVR), as it is the predominant paradigm for enhancing the reasoning capabilities of dLLMs. to elicit the reasoning capability of dLLMs. In this paper, we reveal a counter-intuitive reality: arbitrary order generation, in its current form, narrows rather than expands the reasoning boundary of dLLMs. We find that dLLMs tend to exploit this order flexibility to bypass high-uncertainty tokens that are crucial for exploration, leading to a premature collapse of the solution space. This observation challenges the premise of existing RL approaches for dLLMs, where considerable complexities, such as handling combinatorial trajectories and intractable likelihoods, are often devoted to preserving this flexibility. We demonstrate that effective reasoning is better elicited by intentionally forgoing arbitrary order and applying standard Group Relative Policy Optimization (GRPO) instead. Our approach, JustGRPO, is minimalist yet surprisingly effective (e.g., 89.1% accuracy on GSM8K) while fully retaining the parallel decoding ability of dLLMs. Project page: https://nzl-thu.github.io/the-flexibility-trap
Figure 1: Less flexibility unlocks better reasoning potential. Left: We observe a counter-intuitive phenomenon where restricting dLLMs to standard Autoregressive (AR) order expands the reasoning solution space. Right: Motivated by this, we propose “JustGRPO”. By foregoing complex arbitrary-order adaptations and adopting standard GRPO, we effectively elicit the reasoning capability of dLLMs.
1 Introduction
Recent research has witnessed a surge in Diffusion Large Language Models (dLLMs) Nie et al. (2025); Ye et al. (2025); Zhu et al. (2025); Zhao et al. (2025), which challenge the dominant autoregressive (AR) paradigm Brown et al. (2020); Achiam et al. (2023) by treating sequence generation as a discrete denoising process. Central to the appeal of dLLMs is their theoretical flexibility, which offers two distinct advantages over the strict left-to-right causal chain of AR models: efficient parallel decoding and the capability for arbitrary-order generation.
While the efficiency gains of parallel decoding have been well-established Wu et al. (2025b); Labs et al. (2025); DeepMind (2025); Song et al. (2025); Wu et al. (2025a), the implications of arbitrary-order generation remain less explored Ye et al. (2024); Kim et al. (2025). Theoretically, the unconstrained generation order constitutes a superset of the fixed autoregressive trajectory. This flexibility naturally suggests a potential for superior reasoning: in general reasoning tasks like mathematics and coding, such freedom could unlock non-sequential problem-solving paths inaccessible to standard left-to-right models. As a result, recent works have increasingly adopted RL to elicit reasoning capabilities of dLLMs Zhao et al. (2025); Gong et al. (2025); Wang et al. (2025a); Ou et al. (2025).
In this paper, we present a counter-intuitive observation: arbitrary-order generation, in its current form, narrows rather than expands the reasoning potential elicitable by RL. To rigorously assess this, we employ Pass@$k$ Chen (2021), which measures the coverage of solution space. Recent studies suggest that RL primarily acts to sharpen the base model’s distribution; consequently, the Pass@$k$ performance of the base model effectively sets the upper bound for the reasoning capability of the model after RL training Yue et al. (2025); Liu et al. (2025); Zhang et al. (2025). Under this metric, we compare the reasoning potential of LLaDA Nie et al. (2025) with arbitrary-order generation against standard AR decoding. As shown in Figure 1 (Left), restricting a dLLM to standard AR order in fact yields a higher Pass@$k$, and consequently a higher reasoning boundary, than its flexible counterpart.
Figure 2: Confronting vs. bypassing uncertainty. (a) AR order preserves reasoning space by forcing decisions at uncertain tokens. (b) Arbitrary order bypasses uncertainty and resolves easier tokens first. Once future context is established, the original forks collapse, prematurely narrowing the solution space.
We attribute this counter-intuitive phenomenon to the way model handles uncertainty. Reasoning is inherently non-uniform: it typically hinges on sparse “forking tokens”, i.e., connectives like “Therefore” or “Since” which do not merely continue a sentence but fundamentally steer the logical trajectory into distinct branches Wang et al. (2025c); Cheng et al. (2025); Huang et al. (2025a). At these forks, the reasoning path diverges, naturally manifesting as localized spikes in entropy Wang et al. (2025c). Standard AR decoding compels the model to confront this uncertainty (Figure 2a). By sampling exactly at the fork, the model is able to explore different reasoning paths, thereby preserving the diversity of the generated rationales. Arbitrary-order generation, however, allows the model to bypass these hard decisions (Figure 2b). It prioritizes low-entropy completions first. By the time the model returns to infill the bypassed forks, the established bidirectional context has already severely constrained the potential branches. The ambiguity is prematurely resolved, and the original high entropy is suppressed. We term this phenomenon entropy degradation. Effectively, the model trades the exploration of diverse reasoning paths for the greedy optimization of local consistency.
The above observations motivate a rethink of RL for dLLMs. Current methods operate under the assumption that preserving arbitrary-order flexibility is essential. This commitment incurs a heavy tax: algorithms must grapple with a combinatorial explosion of denoising trajectories Zhao et al. (2025); Yang et al. (2025); Gong et al. (2025) and intractable marginal likelihoods Ou et al. (2025), forcing reliance on unstable approximations Wang et al. (2025a); Ou et al. (2025); Rojas et al. (2025). However, if arbitrary order is non-essential, or even detrimental for eliciting reasoning potential, this complexity is unjustified.
To this end, we propose a return to simplicity with Just GRPO. We demonstrate that eliciting reasoning potential for general reasoning tasks does not require complex, diffusion-specific RL adaptations. Instead, it is best achieved by simply treating the dLLM as an AR model during RL training. This allows us to apply standard GRPO Shao et al. (2024) without bells and whistles, turning an otherwise intractable optimization with unstable approximations into a well-defined task.
JustGRPO is surprisingly effective. On complex reasoning benchmarks, it achieves competitive results (e.g., 89.1% accuracy on GSM8K, 45.1% on MATH), surpassing methods that rely on complex diffusion-specific RL. Crucially, while we train with AR constraints to maximize reasoning potential, the model retains dLLMs’ ability of efficient parallel decoding at inference time. By returning to basic left-to-right ordering, this work encourages a re-evaluation of arbitrary versus AR order in the development of next-generation language models.
2 Preliminaries
2.1 Diffusion Large Language Models
Diffusion Large Language Models (dLLMs), particularly Masked Diffusion Models (MDMs), generate sequences by iteratively denoising a masked state $x_{t}$ initialized from fully masked tokens. The process is indexed by a continuous time variable $t\in[0,1]$, representing the masking ratio. Given a clean sequence $x_{0}$, the forward process independently masks each token with probability $t$:
$q(x_{t}^{k}\mid x_{0}^{k})=\begin{cases}\texttt{[MASK]},&\text{with prob }t,\\ x_{0}^{k},&\text{with prob }1-t.\end{cases}$
Unlike Gaussian diffusion, MDMs directly predict the clean token. A neural network $p_{\theta}(x_{0}\mid x_{t})$ estimates the original token distribution at masked positions. During inference, generation starts from $x_{1}$ (all [MASK]) and iteratively unmasks a subset of tokens based on heuristics such as confidence scores, updating $x_{t}\rightarrow x_{t-\Delta t}$ until completion. As a special case, autoregressive generation can be recovered by always unmasking the leftmost token. The model is trained by minimizing the Negative Evidence Lower Bound, which reduces to a weighted cross entropy loss over masked tokens:
$\mathcal{L}_{\mathrm{MDM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}[0,1],\,x_{t}\sim q(x_{t}\mid x_{0})}\left[\frac{1}{t}\sum_{k=1}^{L}\mathbf{1}[x_{t}^{k}=\texttt{[MASK]}]\,\log p_{\theta}(x_{0}^{k}\mid x_{t})\right].$
2.2 Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO) is a reinforcement learning algorithm that avoids value function estimation by using group level reward normalization. It is typically applied to autoregressive policies $\pi_{\theta}(o_{k}\mid o_{<k},q)$. For each query $q$, GRPO samples a group of $G$ outputs $\{o_{1},\dots,o_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$. An advantage $A_{i}$ is computed by standardizing the reward $r(o_{i})$ against the group statistics: $A_{i}=(r(o_{i})-\mu_{G})/\sigma_{G}$, where $\mu_{G}$ and $\sigma_{G}$ are the group mean and standard deviation. The GRPO objective maximizes a clipped surrogate function with a KL regularization term:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{L_{i}}\sum_{k=1}^{L_{i}}\left(\min\left(\rho_{i,k}A_{i},\text{clip}(\rho_{i,k},1-\epsilon,1+\epsilon)A_{i}\right)-\beta D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\right)\right]$ (1)
with the token level importance ratio $\rho_{i,k}=\frac{\pi_{\theta}(o_{i,k}\mid o_{i,<k},q)}{\pi_{\theta_{\text{old}}}(o_{i,k}\mid o_{i,<k},q)}$.
2.3 Pass@$k$ as a Proxy for Reasoning Potential
To rigorously quantify the reasoning capability boundaries of dLLMs, we adopt the Pass@$k$ metric Chen (2021); Yue et al. (2025). In the context of Reinforcement Learning with Verifiable Rewards (RLVR), which is currently the dominant paradigm for enhancing reasoning capabilities, exploration is a prerequisite for improvement. An RL agent can only reinforce correct reasoning paths if it is capable of sampling them during the exploration phase to obtain a positive reward signal.
Accordingly, Pass@$k$ has been widely established as a standard measure of a model’s reasoning potential Yue et al. (2025); Liu et al. (2025); Zhang et al. (2025). It measures the probability that at least one correct solution is generated within $k$ independent samples, effectively delineating the upper bound of the solution space accessible to the RL optimizer. Formally, following the unbiased estimator formulation Chen (2021); Yue et al. (2025), given $n$ samples where $c$ are correct, Pass@$k$ is calculated as:
$\text{Pass@}k=\mathbb{E}\left[1-\frac{\binom{n-c}{k}}{\binom{n}{k}}\right$] (2)
A high Pass@$k$ indicates that the correct reasoning trajectory lies within the model’s sampling distribution and is thus learnable via RL optimization. Conversely, if a model consistently fails to yield a solution despite a vast sampling budget, it suggests the problem lies beyond its intrinsic reasoning boundary. In such scenarios, standard RLVR methodologies are fundamentally limited by the absence of positive exploration signals Yue et al. (2025).
Figure 3: Reasoning potential measured by Pass@$k$. While arbitrary order is competitive in single-shot settings ($k=1$), it exhibits notably flatter scaling curves compared to AR Order.
3 The Flexibility Trap
In this section, we rigorously test whether the theoretical flexibility of arbitrary-order generation translates into higher reasoning potential. We compare two decoding modes: Arbitrary Order, which follows standard diffusion decoding with low-confidence remasking Nie et al. (2025); Zhao et al. (2025); Wu et al. (2025b), and AR Order, where arbitrary-order flexibility is disabled and generation is strictly constrained to left-to-right decoding. We adopt the commonly used experimental setup in prior diffusion LLM work: a maximum of 256 tokens decoded over 256 steps, a semi-autoregressive block size of 32. Sampling temperature is set to 0.6 as in Yue et al. (2025). The prompt template follows Zhao et al. (2025). Results under alternative temperatures and sampling configurations are deferred to Appendix B.
3.1 Arbitrary Order Limits Reasoning Potential
Pass@$k$ analysis.
We first evaluate the reasoning potential using the Pass@$k$ metric on three representative dLLMs: LLaDA-Instruct Nie et al. (2025), Dream-Instruct Ye et al. (2025), and LLaDA 1.5 Zhu et al. (2025) on four reasoning benchmarks: GSM8K, MATH500, HumanEval, and MBPP. As shown in Figure 3, while arbitrary order often achieves competitive performance at $k=1$, it exhibits a notably flatter scaling curve compared to the AR mode. As $k$ increases, the AR mode demonstrates a stronger capacity to uncover correct solutions.
Figure 4: Solution space coverage measured by Pass@$1024$. The reasoning traces generated by arbitrary order are largely a strict subset of those generated by AR Order.
Solution space coverage.
One might hypothesize that arbitrary order explores a different solution space, albeit less efficiently, which could account for its lower reasoning potential. We test this by analyzing solution coverage at $k=1024$ using LLaDA-Instruct. Figure 4 presents a stark reality: the reasoning traces generated by arbitrary order are largely a strict subset of those generated by AR. On HumanEval, AR solves 21.3% of the problems that arbitrary order misses, whereas the reverse is negligible (0.6%). This indicates that the flexible decoding process rarely unlocks genuinely new solutions. Instead, it appears to retreat into a more conservative subset of the AR solution space.
3.2 Mechanism: The Entropy Degradation
Figure 5: Frequently bypassed tokens in arbitrary order, measured on MATH-500, are typically logical connectors and transition words.
Adaptive decoding bypasses logical forks.
To understand why the theoretically superior solution space of dLLMs collapses in practice, we examine more closely how the two decoding modes handle uncertainty. In AR order, the model is constrained to strictly resolve the left-most unknown token at each step, forcing the model to confront uncertainty as it arises. In contrast, arbitrary order adaptively selects tokens to update based on model confidence, preferentially generating “easy” tokens with high certainty while bypassing “hard” ones. Inspecting the frequently bypassed tokens reveals a clear pattern: As shown in Figure 5, the diffusion sampler disproportionately defers logical connectives and transition markers such as “Therefore”, “Thus”, and “Since”. Prior work has shown that such tokens often have high entropy (which also holds true in dLLMs; see Figure 6), and act as “reasoning sparks” or “logical forks”, functioning as branching points that determine subsequent reasoning directions Wang et al. (2025c); Cheng et al. (2025); Wang et al. (2025b); Huang et al. (2025a). In conventional language models, keeping these tokens in high-entropy state is critical for effective exploration of the reasoning space Wang et al. (2025c).
Figure 6: Entropy degradation. While the global average entropy of Arbitrary Order remains comparable to AR (dashed lines), the entropy at logical forks drops significantly (blue bars).
The “entropy degradation” phenomenon.
This adaptive behavior comes at a cost: the premature collapse of reasoning possibilities. We measure the entropy of these pivotal connectors at decoding. In AR order, these tokens maintain high entropy, reflecting a genuine branching point where multiple logical paths remain viable. In contrast, arbitrary order exhibits a sharp decrease in entropy (Figure 6, with more results in Appendix B). By deferring the logical connector, the model commits to a specific future outcome that is generated based on its inherent inductive biases before deciding the logic that leads there. When the model eventually returns to fill in the bypassed connector, it is no longer making a navigational decision at a fork; it is more like selecting the connector that grammatically bridges the gap to its pre-generated conclusion. The decoding process thus implicitly shifts from reasoning exploration into semantic pattern matching. We term this phenomenon entropy degradation.
Conclusion.
In summary, the flexibility of arbitrary order serves as a mechanism for inference-time exploitation rather than reasoning exploration. By bypassing high-uncertainty tokens, the model effectively collapses the solution space to a safe, low-entropy path, squeezing out slightly better single-shot coherence at the expense of reasoning potential. Autoregressive models, by contrast, lack this bypassing capability and are therefore forced to sample directly from high-entropy distributions at logical forks. It is precisely this inability to circumvent critical decision points that prevents the premature narrowing of the search space and preserves the reasoning potential.
4 “Just GRPO” for Diffusion Language Models
The findings in Section 3 suggest that arbitrary order actually limits the reasoning potential accessible to RL. Despite this, current RL methods for dLLMs remain heavily burdened by the need to preserve this specific flexibility. In this section, we uncover the heavy “tax” imposed by this flexibility (Section 4.1) and show that discarding it enables a minimalist yet surprisingly effective solution: JustGRPO (Section 4.2).
4.1 The Flexibility Tax in dLLMs’ RL
Existing diffusion RL methods operate under the premise that the policy must optimize over the full combinatorial space of denoising trajectories $\mathcal{T}$ to preserve the flexibility of arbitrary order generation. This design choice, while conceptually general, introduces several fundamental challenges.
Ambiguity in token-level decomposition.
In dLLMs, a generation state $s_{t}$ is a noisy sequence conditioned on a stochastic unmasking trajectory $\tau$. Unlike autoregressive models, dLLMs do not admit a unique, index-aligned conditional probability of the form $\pi(o_{t}\mid s_{t})$, making token-level credit assignment ambiguous and rendering the standard importance ratio $\rho_{t}=\frac{\pi_{\theta}(o_{t}\mid s_{t})}{\pi_{\text{old}}(o_{t}\mid s_{t})}$ hard to define.
Intractable sequence likelihood.
Autoregressive models factorize the sequence likelihood as $\log\pi(o)=\sum_{t}\log\pi(o_{t}\mid o_{<t},q)$, whereas dLLMs require marginalization over all valid denoising trajectories, $\pi_{\theta}(o\mid q)=\sum_{\tau\in\mathcal{T}}\pi_{\theta}(o,\tau\mid q)$. For a sequence of length $N$, the trajectory space grows as $|\mathcal{T}|=O(N!)$, rendering exact likelihood computation infeasible and forcing existing methods to rely on ELBO-based surrogates rather than the true objective Wang et al. (2025a); Ou et al. (2025).
Sampler-learner mismatch.
Even with an accurate likelihood approximation, a more subtle issue persists. In practice, rollout samples are produced by heuristic-guided policies $o\sim\pi_{\theta}^{\text{heur}}(o\mid q)$ to explore the combinatorial space. However, the ELBO objective still targets the likelihood of the original model distribution $\pi_{\theta}(o\mid q)$, rather than that of the heuristic-guided sampler $\pi_{\theta}^{\text{heur}}(o\mid q)$, leading to a critical mismatch between sampling and optimization that can degrade performance Schulman et al. (2015).
4.2 JustGRPO
We propose a return to simplicity. Since pure autoregressive order yields better reasoning potential (Section 3), we explicitly forgo arbitrary-order generation during the RL stage. This constraint transforms the dLLM from a chaotic sequence denoiser into a well-defined autoregressive policy $\pi_{\theta}^{AR}$.
Formulation.
Standard GRPO assumes a policy $\pi(o_{t}|o_{<t},q)$ accepting a partial sequence $o_{<t}$ with all tokens observed and predicts one token $o_{t}$ at a time, where $q$ is the query. Diffusion language models, however, are architected as sequence-level denoisers that accept a full sequence with mixed observed and masked tokens and predict the original values for all masked positions simultaneously.
By forgoing arbitrary-order generation, we are able to bridge the above gap and rigorously define an AR policy $\pi_{\theta}^{\text{AR}}$ for dLLMs. To obtain the probability of the next token $o_{t}$ given history $o_{<t}$, we construct an input state $\tilde{x}_{t}$ where the past is observed and the future is masked:
$\tilde{x}_{t}=[\underbrace{o_{1},\dots,o_{t-1}}_{\text{Observed}},\underbrace{\texttt{[MASK]},\dots,\texttt{[MASK]}}_{\text{Masked}}].$ (3)
Although the diffusion language model outputs predictions for all masked positions, the autoregressive policy concerns only the next token $o_{t}$. We thus define $\pi_{\theta}^{\text{AR}}(o_{t}|o_{<t},q)$ as the probability distribution of $o_{t}$:
$\pi_{\theta}^{\text{AR}}(o_{t}|o_{<t},q)\triangleq\text{Softmax}(f_{\theta}(\tilde{x}_{t}))_{t},$ (4)
where $f_{\theta}$ denotes the model logits. Consequently, the likelihood of a complete reasoning chain $o$ is exactly computable as:
$\pi_{\theta}^{\text{AR}}(o|q)=\prod_{t=1}^{L}\pi_{\theta}^{\text{AR}}(o_{t}|o_{<t},q).$ (5)
Optimization.
The above formulation enables the direct application of standard GRPO to diffusion language models. For each query $q$, we sample a group of outputs $\{o^{i}\}_{i=1}^{G}$ using the old policy $\pi_{\theta_{\text{old}}}^{\text{AR}}$. The objective is:
$\mathcal{J}(\theta)$ $=\mathbb{E}_{q\sim P(Q),\{o^{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}^{\text{AR}}}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o^{i}|}\sum_{t=1}^{|o^{i}|}\left(\min\left(\rho_{i,t}\hat{A}_{i,t},\text{clip}(\rho_{i,t},1-\varepsilon,1+\varepsilon)\hat{A}_{i,t}\right)-\beta\mathbb{D}_{\text{KL}}\right)\right],$ (6)
where $\rho_{i,t}=\frac{\pi_{\theta}^{\text{AR}}(o^{i}_{t}\mid o^{i}_{<t},q)}{\pi_{\theta_{\text{old}}}^{\text{AR}}(o^{i}_{t}\mid o^{i}_{<t},q)}$ is the probability ratio between the current and old policies and $\hat{A}_{i,t}$ denotes the group-standardized advantage.
Remarks.
One might worry whether training in AR mode turns the diffusion language model into a standard autoregressive model. This is not the case. The AR constraint is applied only during training to correctly assign credit. It refines the model’s joint distribution $p(o)$ without altering the underlying architecture. At inference time, the model retains its conditional independence properties, still allowing us to employ parallel samplers Ben-Hamu et al. (2025) to accelerate decoding. JustGRPO thus achieves the reasoning depth of autoregressive models while preserving the inference speed of dLLMs (see Section 5.2).
5 Experiments
We evaluate JustGRPO on standard mathematical reasoning and coding benchmarks. Our experimental design aims to verify two hypotheses: (i) that enforcing autoregressive (AR) order during RL training elicits superior reasoning capabilities compared to complex arbitrary-order approximations, and (ii) that this constraint applies only to the optimization objective, leaving the diffusion model’s parallel decoding capabilities intact at inference.
Experimental Setups.
We apply JustGRPO on LLaDA-Instruct Nie et al. (2025) and evaluate its effectiveness on four standard reasoning and coding benchmarks: GSM8K, MATH, HumanEval, and MBPP. For mathematical tasks, we train on the official training split of each dataset following Zhao et al. (2025); Ou et al. (2025). For coding tasks, we train on a subset of AceCoder-87K Zeng et al. (2025) following Gong et al. (2025); Ou et al. (2025). We apply JustGRPO directly to the models without additional task-specific SFT. Our training recipe largely follows Huang et al. (2025b). To evaluate the trained models, we follow the standard LLaDA evaluation protocol Nie et al. (2025), which applies low-confidence remasking together with semi-autoregressive decoding in blocks of 32 tokens, using a sampling temperature of 0. We evaluate all benchmarks at generation lengths of 128, 256, and 512 following Zhao et al. (2025); Ou et al. (2025). More details on training are provided in Appendix A.
5.1 Main Results
Table 1 reports the system-level comparison. We observe that simplifying the training objective to a standard autoregressive formulation yields consistent improvements over methods specifically designed for dLLMs.
Performance on reasoning tasks.
On GSM8K, JustGRPO achieves 89.1% accuracy (seq len 256), improving upon the previous best method, SPG, by a non-trivial margin (3.0%). This trend generalizes to the more challenging MATH-500 benchmark, where our approach outperforms ESPO by 6.1%. These results challenge the prevailing assumption that RL for diffusion models requires optimizing over the full combinatorial space of denoising trajectories. Instead, treating the dLLM as a sequential generator during training appears more effective for credit assignment in logic-heavy tasks.
Robustness across generation budgets.
A potential concern is that AR constraints might overfit the model to specific trajectory lengths. However, we observe robust performance gains across varying sequence lengths (128, 256, 512). This stability suggests that the policy has improved its fundamental reasoning capability, i.e., the ability to navigate logical branches, rather than merely memorizing fixed-length patterns.
Table 1: System-level comparison. RL post-training approaches on LLaDA-Instruct. JustGRPO consistently achieves state-of-the-art performance across all tasks and sequence lengths. LLaDA-1.5 and LLADOU are excluded from the comparison as LLaDA-1.5 is trained on a privately collected dataset at a significantly larger scale, while LLADOU modifies the base architecture with an auxiliary module.
GSM8K MATH-500 HumanEval MBPP
Model / Seq Len 128 256 512 128 256 512 128 256 512 128 256 512
LLaDA-1.5 - 83.3 - - - - 29.3 39.6 51.9 39.6 39.9 38.8
LLaDOU - 88.1 - - 41.1 - - 59.1 - - 51.6 -
D1 73.2 81.1 82.1 33.8 38.6 40.2 - - - - - -
WD1 77.2 80.8 82.3 33.3 34.4 39.0 - - - - - -
$d$-TreeRPO - 81.2 82.6 - 37.7 38.9 - - - - - -
ESPO 80.0 82.3 83.7 36.0 39.0 43.4 28.1 42.1 50.0 47.4 44.6 44.2
GDPO 78.4 82.8 84.5 33.2 39.6 41.4 26.2 39.6 39.0 43.6 50.6 47.1
SPG 78.5 86.1 84.5 33.4 40.0 41.8 - - - - - -
JustGRPO (Ours) 83.8 89.1 89.8 39.0 45.1 45.2 37.8 49.4 48.7 50.6 52.4 49.0
5.2 JustGRPO Preserves Parallel Decoding
Figure 7: JustGRPO preserves the parallel decoding capability of dLLMs. Interestingly, when compared to original instruct model, accuracy gains are larger with more parallel tokens, likely due to more robust reasoning capabilities after JustGRPO training. We adopt training-free EB-sampler Ben-Hamu et al. (2025) for parallel decoding.
We further investigate whether the AR training constraint compromises the model’s inherent diffusion capabilities. We employ the training-free Entropy Bounded (EB) Sampler Ben-Hamu et al. (2025) to evaluate inference performance under varying degrees of parallelism (tokens per step).
Figure 7 reveals that our model not only retains full compatibility with parallel decoding but exhibits a strictly superior trade-off between speed and accuracy. Surprisingly, the performance gain becomes more pronounced as parallelism increases. As shown in the MBPP and HumanEval results, while the baseline’s performance degrades sharply with more aggressive parallel steps, our model maintains stability. For instance, on MBPP, the accuracy gap expands from +10.6% at conservative settings (1 token/step) to +25.5% at aggressive settings ($\sim$5 tokens/step).
This observation suggests that JustGRPO does not merely fit a specific decoding path; rather, it learns a more robust reasoning manifold that is resilient to the approximation errors inherent in parallel sampling. The AR formulation thus acts as an effective training scaffold that refines the joint distribution $p(o)$ during optimization, providing a more stable foundation for parallel samplers to exploit at inference.
6 Related Work
Diffusion language models.
Inspired by the success of diffusion models in continuous image domains Ho et al. (2020); Rombach et al. (2022), recent work has extended diffusion to discrete text generation. Early approaches operating in continuous embedding spaces Li et al. (2022); Gong et al. (2022); Han et al. (2022) suffered from optimization and discretization issues. In contrast, masked diffusion models Lou et al. (2023); Sahoo et al. (2024); Shi et al. (2024); Ou et al. (2024), which define the diffusion process directly in the discrete token space via random masking, have become the dominant paradigm. Notably, recent large-scale models such as LLaDA Nie et al. (2025) and Dream Ye et al. (2025) demonstrate scalability and performance competitive with autoregressive (AR) models. The emergence of such large-scale diffusion language models has spurred a growing body of follow-up studies exploring their distinctive modeling and inference properties. Among various advantages discussed in the literature, two aspects have received particular attention. First, dLLMs naturally support parallel decoding, enabling significant inference acceleration Wu et al. (2025b); Ben-Hamu et al. (2025); Labs et al. (2025); DeepMind (2025); Song et al. (2025). Second, their non-autoregressive formulation allows for arbitrary-order token generation, which has been hypothesized to benefit complex reasoning by relaxing strict left-to-right constraints Ye et al. (2024); Kim et al. (2025).
The value of order arbitrariness.
While early studies validated the utility of arbitrary order generation in constrained tasks like Sudoku and Zebra Puzzles Ye et al. (2024); Nie et al. (2025); Kim et al. (2025), recent research has begun to investigate its value in standard reasoning domains. One line of research, exemplified by Dream-Coder Xie et al. (2025) and DiffuCoder Gong et al. (2025), shows that diffusion models can naturally exhibit non-standard, human-like decoding behaviors (e.g., sketch-first reasoning) without any explicit supervision on generation order. DiffuCoder Gong et al. (2025) further quantifies this phenomenon by demonstrating that higher sampling temperatures reduce AR-ness, with increased order randomness correlating with improved output diversity. Another line of research, exemplified by P2 Peng et al. (2025) and LLaDOU Huang et al. (2025b), explicitly optimizes the decoding strategy for better generation order, thus better performance. Despite these advances, it remains unclear whether the observed gains primarily arise from better exploitation of existing solution patterns encoded in the data and model, or whether order arbitrariness itself enables qualitatively new reasoning strategies that are unattainable under a purely autoregressive decoding regime.
Reinforcement learning for diffusion language models.
Reinforcement learning for dLLMs faces structural optimization hurdles distinct from the autoregressive paradigm, primarily stemming from the combinatorial explosion of denoising trajectories. While early attempts Zhao et al. (2025); Yang et al. (2025); Gong et al. (2025) sought to adapt token-level formulations directly, they were fundamentally limited by ill-defined state transitions, necessitating reliance on unstable mean-field approximations. Consequently, the field has shifted toward sequence-level perspectives Zhu et al. (2025); Wang et al. (2025a); Rojas et al. (2025); Ou et al. (2025), employing various surrogates to approximate the intractable marginal likelihood. Yet, a critical off-policy misalignment persists across these methods: the heuristic-guided sampling required for efficient exploration diverges from the underlying diffusion prior, rendering gradients biased without principled correction Schulman et al. (2015). Two notable exceptions partially address these issues. LLaDOU Huang et al. (2025b) explicitly models token position selection via an auxiliary policy, enabling direct estimation of trajectory likelihoods, while TraceRL Wang et al. (2025d) aligns optimization with inference traces through a shrinkage-based step-wise MDP formulation. Nevertheless, these approaches remain committed to preserving the full arbitrary-order generation mechanism, implicitly treating its structural complexity as indispensable, whereas we question whether effective RL training can be achieved more simply by reexamining the necessity of arbitrary order itself.
7 Conclusion
The intuitive appeal of diffusion language models (dLLMs) lies in their order arbitrariness, often perceived as a superior mechanism for navigating complex reasoning paths. Our study reveals a counter-intuitive reality: this unrestricted flexibility in fact narrows the reasoning potential. By allowing the model to bypass high-entropy tokens, effectively skipping the most demanding logical branches, arbitrary-order generation acts as an exploitation mechanism that prioritizes greedy optimization of individual trajectories at the expense of broader solution coverage.
Therefore, eliciting the reasoning capability of dLLMs can be simpler. By operating dLLMs in a standard autoregressive manner, we enable the direct application of Group Relative Policy Optimization (GRPO) without any complex adaptations tailored for order arbitrariness. This intentional constraint paradoxically yields a significant upgrade in reasoning performance, while fully preserving the parallel decoding capabilities of dLLMs. By returning to the basic, natural left-to-right order of language modeling, we hope to encourage a re-examination of its real value in the training of next-generation diffusion models.
Appendix
Appendix A Experimental Details
A.1 Data Preparation
For mathematical reasoning tasks, we train on the official training split of each dataset, following the standard protocol in Zhao et al. (2025); Ou et al. (2025). For code generation tasks, we adopt the AceCoder-87K dataset Zeng et al. (2025). Following the data processing pipeline of DiffuCoder Gong et al. (2025), we further select 21K challenging samples from AceCoder-87K that are equipped with verifiable unit tests.
A.2 Training Configuration
[…]
Table 2: Training hyperparameters for JustGRPO.
Hyperparameter Value
Base Model LLaDA 8B Instruct
RL Algorithm GRPO
Optimizer AdamW
Learning Rate $5\times 10^{-6}$
LR Scheduler Constant
Weight Decay 0.0
Optimizer Betas $(\beta_{1},\beta_{2})$ $(0.9,0.999)$
Global Batch Size 64
Group Size ($G$) 16
Policy Update Steps 1
Training Steps 125
Max Completion Length 256
Sampling Temperature 1.0
KL Penalty Coefficient 0.0
A.3 Reward Function
Mathematical Reasoning Tasks.
We employ a binary reward scheme. Each completion receives a reward of 1 if and only if the final answer is mathematically equivalent to the ground-truth solution, and 0 otherwise, following Huang et al. (2025b).
Code Generation Tasks.
For code generation, the total reward $r$ is defined as a weighted sum of a correctness reward $r_{\text{code}}$ and a format reward $r_{\text{format}}$:
$r=r_{\text{code}}+r_{\text{format}}.$ (7)
  • Correctness reward ($r_{\text{code}}$): Defined as the pass rate (ranging from 0 to 1) of the generated code on the provided unit tests. This term is only evaluated when $r_{\text{format}}=1$.
  • Format reward ($r_{\text{format}}$): A heuristic reward designed to encourage syntactically valid outputs.
    • $1.0$: A valid Markdown code block with correct Python syntax.
    • $0.5$: A valid Markdown code block that contains syntax errors.
    • $0.0$: Failure to generate a valid Markdown code block.
Appendix B More Analysis Results
To further validate the robustness of our findings in Section 3, we conduct extended analyses on the HumanEval benchmark using the LLaDA-Instruct model Nie et al. (2025). We investigate the impact of temperature, sampling strategies, and decoding block sizes on reasoning performance.
B.1 Temperature analysis.
Figure 8: Pass@K comparison with different temperatures.
As shown in Figure˜8, the AR mode exhibits a standard pattern: performance peaks at moderate temperatures ($T\approx 0.6$) and degrades when $T>1.0$, whereas the Arbitrary Order mode attains its peak performance at higher temperatures.
This observation aligns with the entropy degradation mechanism discussed in Section 3: the diffusion sampler inherently suppresses uncertainty at critical branching points, thereby requiring higher temperatures to induce sufficient exploration. Crucially, even under these optimized settings, the arbitrary order mode fails to match the reasoning potential of the AR mode. As illustrated in the “Optimal” comparison setting (bottom right of Figure˜8), the best-performing AR configuration still outperforms the optimal arbitrary order baseline in scaling behavior. A plausible explanation is that excessively high temperatures in arbitrary order decoding inject noise into tokens that require high determinism (e.g., code syntax or mathematical suffixes), leading to nonsensical outputs and degraded results Wang et al. (2025c).
B.2 Different sampling algorithms.
We also experiment with different sampling algorithms, such as negative entropy sampling (Neg-Entropy)  Ye et al. (2025) and top-$k$ margin sampling (Margin)  Kim et al. (2025) in Figure 9(a).
Figure 9: (a) Sampling algorithm comparison. (b) Correlation between different sampling algorithms and AR in per-problem accuracy. We compare different sampling algorithms in pass@k performance and correlation with AR in per-problem accuracy.
The results show that although more sophisticated sampling algorithms can achieve better pass@k performance than default confidence-based sampling, they still cannot catch up with AR order. Meanwhile, these better pass@k algorithms also show slightly worse pass@1 performance compared to confidence-based sampling, making the overall pass@$k$ curve closer to AR order’s pass@$k$ curve. To investigate this similarity, we calculate the per-problem accuracy correlation between different sampling algorithms and the AR mode (Figure 9(b)). We observe that algorithms with higher scaling potential (higher pass@128) consistently show stronger correlation with AR, with the most effective method (Neg-Entropy) achieving a correlation of $0.970$. This suggests that sampling algorithms with better pass@$k$ tend to behave more like AR in terms of task-level performance characteristics.
B.3 Block size analysis.
Figure 10: Pass@$k$ comparison with different semi-autoregressive block sizes of arbitrary order (AO) generation. Smaller block sizes explicitly restrict the model’s order flexibility, leading to a more AR-like behavior and better pass@k performance.
We analyze the effect of the semi-autoregressive block size in Figure˜10. The AR Order (which is effectively block size of 1) maintains a clear advantage across all tested block sizes. Furthermore, we observe a trend of improvement in pass@$k$ as the block size decreases (from 128 to 8). Since smaller block sizes explicitly restrict the model’s order flexibility, this trend is aligned with our finding that less order flexibility leads to better reasoning potential.
B.4 Entropy Comparison Results on More Forking Tokens
Figure 11: Entropy comparison results on more forking tokens.
To validate the robustness of the “entropy degradation” phenomenon observed in Section 3, we extended our analysis to a wider range of logical connectors. We conducted experiments on a comprehensive set of common logical connectors that typically serve as pivotal decision points in reasoning chains.
The results, shown in Figure˜11, demonstrate that the phenomenon is consistent across these diverse tokens. Similar to the primary findings, the AR order maintains higher average entropy at these forks, indicating active reasoning and decision-making. Conversely, the Arbitrary Order consistently results in lower entropy, confirming the premature collapse of branching possibilities.
The specific forking tokens evaluated in this extended experiment include: “Therefore”, “Thus”, “So”, “Since”, “When”, “Given”, “However”, “Let”, “First”, “Then”, “Next”, “Finally”, “Now”, “Similarly”, “Calculate”, “Solving”, “Notice”, “Specifically”, “Follows”, “Because”, “But”, “Or”, “Consider”, “Also”, “Express”, and “Write”.
B.5 Training Efficiency Analysis
[…]
Applying GRPO to dLLMs presents a structural trade-off. Unlike autoregressive models that benefit from causal masking, dLLMs require independent likelihood evaluation at each position, inherently incurring additional computational overhead. While approximation-based methods Zhao et al. (2025); Ou et al. (2025) bypass this cost by forgoing exact estimation, we demonstrate that adhering to exact likelihoods yields a superior trade-off between accuracy and wall-clock time.
Figure 12 compares JustGRPO against ESPO Ou et al. (2025), a representative approximation-based baseline. It can be observed that ESPO saturates early, effectively plateauing around 10 hours. In contrast, JustGRPO exhibits a superior efficiency frontier: it surpasses the baseline’s peak accuracy in less time and maintains a continuous scaling trend. Furthermore, motivated by the observation that reasoning is primarily steered by a subset of high-entropy forking tokens Wang et al. (2025c), we demonstrate that JustGRPO allows for further acceleration via a simple heuristic: restricting gradient computation to the top-25% high-entropy tokens (Ours-Fast). This modification yields even faster convergence without compromising the final performance. We prioritize simplicity in the current implementation and leave further efficiency optimizations for future work.

### v2

@@
 We find that dLLMs tend to exploit this order flexibility to bypass high-uncertainty tokens that are crucial for exploration, leading to a premature collapse of the solution space.
-This observation challenges the premise of existing RL approaches for dLLMs, where considerable complexities, such as handling combinatorial trajectories and intractable likelihoods, are often devoted to preserving this flexibility.
-We demonstrate that effective reasoning is better elicited by intentionally forgoing arbitrary order and applying standard Group Relative Policy Optimization (GRPO) instead.
+This observation motivates a rethink of RL approaches for dLLMs, where considerable complexities, such as handling combinatorial trajectories and intractable likelihoods, are often devoted to preserving this flexibility.
+We demonstrate that effective reasoning can be better elicited by intentionally forgoing arbitrary order and applying standard Group Relative Policy Optimization (GRPO) instead.
 Our approach, JustGRPO, is minimalist yet surprisingly effective (e.g., 89.1% accuracy on GSM8K) while fully retaining the parallel decoding ability of dLLMs.
@@
 Right: Motivated by this, we propose “JustGRPO”.
-By foregoing complex arbitrary-order adaptations and adopting standard GRPO, we effectively elicit the reasoning capability of dLLMs.
+By foregoing complex arbitrary-order adaptations and adopting just GRPO, we elicit the reasoning capability of dLLMs more effectively.
+Results reported with generation length of 256, more results in Table 1.
 1 Introduction
@@
 (2025).
-Theoretically, the unconstrained generation order constitutes a superset of the fixed autoregressive trajectory.
-This flexibility naturally suggests a potential for superior reasoning: in general reasoning tasks like mathematics and coding, such freedom could unlock non-sequential problem-solving paths inaccessible to standard left-to-right models.
+In theory, the unconstrained generation order constitutes a superset of the fixed autoregressive trajectory, unlocking non-sequential problem-solving paths.
+Ideally, such advantage could facilitate more advanced problem-solving abilities that benefit general reasoning tasks like mathematics and coding.
 As a result, recent works have increasingly adopted RL to elicit reasoning capabilities of dLLMs Zhao et al.
@@
 (2025); Wang et al.
-(2025a); Ou et al.
-(2025).
+(2025a).
 In this paper, we present a counter-intuitive observation: arbitrary-order generation, in its current form, narrows rather than expands the reasoning potential elicitable by RL.
@@
 Arbitrary-order generation, however, allows the model to bypass these hard decisions (Figure 2b).
-It prioritizes low-entropy completions first.
-By the time the model returns to infill the bypassed forks, the established bidirectional context has already severely constrained the potential branches.
+It prioritizes low-uncertainty completions first.
+By the time the model returns to infill the bypassed forks, the established future context has already severely constrained the potential branches.
 The ambiguity is prematurely resolved, and the original high entropy is suppressed.
@@
 We demonstrate that eliciting reasoning potential for general reasoning tasks does not require complex, diffusion-specific RL adaptations.
-Instead, it is best achieved by simply treating the dLLM as an AR model during RL training.
+Instead, it can be better achieved by simply treating the dLLM as an AR model during RL training.
 This allows us to apply standard GRPO Shao et al.
@@
 JustGRPO is surprisingly effective.
-On complex reasoning benchmarks, it achieves competitive results (e.g., 89.1% accuracy on GSM8K, 45.1% on MATH), surpassing methods that rely on complex diffusion-specific RL.
+On complex reasoning benchmarks, it achieves competitive results (e.g., 89.1% accuracy on GSM8K, 45.1% on MATH-500), surpassing methods that rely on complex diffusion-specific RL.
 Crucially, while we train with AR constraints to maximize reasoning potential, the model retains dLLMs’ ability of efficient parallel decoding at inference time.
@@
 Given a clean sequence $x_{0}$, the forward process independently masks each token with probability $t$:
-$q(x_{t}^{k}\mid x_{0}^{k})=\begin{cases}\texttt{[MASK]},&\text{with prob }t,\\ x_{0}^{k},&\text{with prob }1-t.\end{cases}$
-Unlike Gaussian diffusion, MDMs directly predict the clean token.
+$q(x_{t,k}\mid x_{0,k})=\begin{cases}\texttt{[MASK]},&\text{with prob }t,\\ x_{0,k},&\text{with prob }1-t,\end{cases}$ (1)
+where $k$ is the token index.
 A neural network $p_{\theta}(x_{0}\mid x_{t})$ estimates the original token distribution at masked positions.
 During inference, generation starts from $x_{1}$ (all [MASK]) and iteratively unmasks a subset of tokens based on heuristics such as confidence scores, updating $x_{t}\rightarrow x_{t-\Delta t}$ until completion.
-As a special case, autoregressive generation can be recovered by always unmasking the leftmost token.
+As a special case, dLLMs can also perform generation autoregressively by always unmasking the leftmost token.
 The model is trained by minimizing the Negative Evidence Lower Bound, which reduces to a weighted cross entropy loss over masked tokens:
-$\mathcal{L}_{\mathrm{MDM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}[0,1],\,x_{t}\sim q(x_{t}\mid x_{0})}\left[\frac{1}{t}\sum_{k=1}^{L}\mathbf{1}[x_{t}^{k}=\texttt{[MASK]}]\,\log p_{\theta}(x_{0}^{k}\mid x_{t})\right].$
+$\mathcal{L}_{\mathrm{MDM}}(\theta)=-\mathbb{E}_{t\sim\mathcal{U}[0,1],\,x_{t}\sim q(x_{t}\mid x_{0})}\left[\frac{1}{t}\sum_{k=1}^{L}\mathbf{1}[x_{t,k}=\texttt{[MASK]}]\,\log p_{\theta}(x_{0,k}\mid x_{t})\right].$ (2)
 2.2 Group Relative Policy Optimization
@@
 For each query $q$, GRPO samples a group of $G$ outputs $\{o_{1},\dots,o_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$.
-An advantage $A_{i}$ is computed by standardizing the reward $r(o_{i})$ against the group statistics: $A_{i}=(r(o_{i})-\mu_{G})/\sigma_{G}$, where $\mu_{G}$ and $\sigma_{G}$ are the group mean and standard deviation.
+An advantage $\hat{A}_{i,k}$ is computed by standardizing the reward $r(o_{i})$ against the group statistics: $\hat{A}_{i,k}=(r(o_{i})-\mu_{G})/\sigma_{G}$, where $\mu_{G}$ and $\sigma_{G}$ are the group mean and standard deviation.
 The GRPO objective maximizes a clipped surrogate function with a KL regularization term:
-$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{L_{i}}\sum_{k=1}^{L_{i}}\left(\min\left(\rho_{i,k}A_{i},\text{clip}(\rho_{i,k},1-\epsilon,1+\epsilon)A_{i}\right)-\beta D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\right)\right]$ (1)
+$\mathcal{J}(\theta)=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\left(\min\left(\rho_{i,k}\hat{A}_{i,k},\text{clip}(\rho_{i,k},1-\epsilon,1+\epsilon)\hat{A}_{i,k}\right)-\beta D_{\text{KL}}\right)\right]$ (3)
 with the token level importance ratio $\rho_{i,k}=\frac{\pi_{\theta}(o_{i,k}\mid o_{i,<k},q)}{\pi_{\theta_{\text{old}}}(o_{i,k}\mid o_{i,<k},q)}$.
@@
 (2025).
-It measures the probability that at least one correct solution is generated within $k$ independent samples, effectively delineating the upper bound of the solution space accessible to the RL optimizer.
+It measures the probability that at least one correct solution is generated within $k$ independent samples, effectively representing the upper bound of the solution space accessible to RL optimization.
 Formally, following the unbiased estimator formulation Chen (2021); Yue et al.
@@
 (2025).
-Figure 3: Reasoning potential measured by Pass@$k$. While arbitrary order is competitive in single-shot settings ($k=1$), it exhibits notably flatter scaling curves compared to AR Order.
+Figure 3: Reasoning potential measured by Pass@$k$. While arbitrary order is competitive in single-shot settings ($k=1$), it exhibits flatter scaling curves compared to AR Order.
 3 The Flexibility Trap
-In this section, we rigorously test whether the theoretical flexibility of arbitrary-order generation translates into higher reasoning potential.
+In this section, we rigorously test whether the flexibility of arbitrary-order generation translates into higher reasoning potential.
 We compare two decoding modes: Arbitrary Order, which follows standard diffusion decoding with low-confidence remasking Nie et al.
@@
 (2025), and LLaDA 1.5 Zhu et al.
-(2025) on four reasoning benchmarks: GSM8K, MATH500, HumanEval, and MBPP.
+(2025) on four reasoning benchmarks: GSM8K, MATH-500, HumanEval, and MBPP.
 As shown in Figure 3, while arbitrary order often achieves competitive performance at $k=1$, it exhibits a notably flatter scaling curve compared to the AR mode.
 As $k$ increases, the AR mode demonstrates a stronger capacity to uncover correct solutions.
-Figure 4: Solution space coverage measured by Pass@$1024$. The reasoning traces generated by arbitrary order are largely a strict subset of those generated by AR Order.
+Figure 4: Solution space coverage measured by Pass@$1024$. The problems solvable by arbitrary order (AO) are largely a subset of those solved by AR Order.
 Solution space coverage.
-One might hypothesize that arbitrary order explores a different solution space, albeit less efficiently, which could account for its lower reasoning potential.
+One might wonder whether arbitrary order explores a different solution space, albeit less efficiently, which could account for its lower reasoning potential.
 We test this by analyzing solution coverage at $k=1024$ using LLaDA-Instruct.
-Figure 4 presents a stark reality: the reasoning traces generated by arbitrary order are largely a strict subset of those generated by AR.
-On HumanEval, AR solves 21.3% of the problems that arbitrary order misses, whereas the reverse is negligible (0.6%).
+Figure 4 shows that the problems solvable by arbitrary order are largely a subset of those solved by AR.
+On HumanEval, 21.3% of problems are solved exclusively by AR, whereas the reverse is negligible (0.6%).
 This indicates that the flexible decoding process rarely unlocks genuinely new solutions.
@@
 Adaptive decoding bypasses logical forks.
-To understand why the theoretically superior solution space of dLLMs collapses in practice, we examine more closely how the two decoding modes handle uncertainty.
+To understand why the theoretically superior solution space of dLLMs collapses in practice, we examine more closely how the two decoding modes behave.
 In AR order, the model is constrained to strictly resolve the left-most unknown token at each step, forcing the model to confront uncertainty as it arises.
@@
 (2025c).
-Figure 6: Entropy degradation. While the global average entropy of Arbitrary Order remains comparable to AR (dashed lines), the entropy at logical forks drops significantly (blue bars).
+Figure 6: Entropy degradation. While the global average token entropy (dashed lines) of arbitrary order remains comparable to AR, the entropy at logical forks (blue bars) drops significantly.
 The “entropy degradation” phenomenon.
-This adaptive behavior comes at a cost: the premature collapse of reasoning possibilities.
-We measure the entropy of these pivotal connectors at decoding.
-In AR order, these tokens maintain high entropy, reflecting a genuine branching point where multiple logical paths remain viable.
-In contrast, arbitrary order exhibits a sharp decrease in entropy (Figure 6, with more results in Appendix B).
-By deferring the logical connector, the model commits to a specific future outcome that is generated based on its inherent inductive biases before deciding the logic that leads there.
-When the model eventually returns to fill in the bypassed connector, it is no longer making a navigational decision at a fork; it is more like selecting the connector that grammatically bridges the gap to its pre-generated conclusion.
-The decoding process thus implicitly shifts from reasoning exploration into semantic pattern matching.
+How does the adaptive behavior of arbitrary order affect these logical forking tokens?
+We measure the entropy of them when they are decoded in Figure 6 (more results in Appendix C).
+In AR order, these tokens generally maintain high entropy when decoded, indicating a relatively fruitful branching point where multiple reasoning paths remain viable Wang et al.
+(2025c); Cheng et al.
+(2025).
+In contrast, arbitrary order exhibits a sharp decrease in entropy.
+By deferring the logical connectors, the model prioritizes generating easier future tokens before deciding the logic connections that lead to them.
+When the model eventually returns to fill in the bypassed connectors, the navigational freedom is notably constrained; the process acts less as an open-ended navigational decision at a fork, and more as a retrospective alignment to bridge the gap to the pre-generated conclusion.
 We term this phenomenon entropy degradation.
@@
 In summary, the flexibility of arbitrary order serves as a mechanism for inference-time exploitation rather than reasoning exploration.
-By bypassing high-uncertainty tokens, the model effectively collapses the solution space to a safe, low-entropy path, squeezing out slightly better single-shot coherence at the expense of reasoning potential.
-Autoregressive models, by contrast, lack this bypassing capability and are therefore forced to sample directly from high-entropy distributions at logical forks.
+By bypassing high-uncertainty tokens, the model effectively collapses the solution space to a safer, lower-entropy path, squeezing out slightly better single-shot coherence at the expense of reasoning potential.
+Autoregressive ordering, by contrast, lack this bypassing capability and are therefore forced to sample directly from high-entropy distributions at logical forks.
 It is precisely this inability to circumvent critical decision points that prevents the premature narrowing of the search space and preserves the reasoning potential.
@@
 4.1 The Flexibility Tax in dLLMs’ RL
-Existing diffusion RL methods operate under the premise that the policy must optimize over the full combinatorial space of denoising trajectories $\mathcal{T}$ to preserve the flexibility of arbitrary order generation.
-This design choice, while conceptually general, introduces several fundamental challenges.
+Existing diffusion RL methods operate under the premise that the policy must accommodate the combinatorial space of denoising trajectories $\mathcal{T}$ to preserve arbitrary order generation.
+This design choice, while conceptually general, introduces several fundamental challenges:
 Ambiguity in token-level decomposition.
@@
 Autoregressive models factorize the sequence likelihood as $\log\pi(o)=\sum_{t}\log\pi(o_{t}\mid o_{<t},q)$, whereas dLLMs require marginalization over all valid denoising trajectories, $\pi_{\theta}(o\mid q)=\sum_{\tau\in\mathcal{T}}\pi_{\theta}(o,\tau\mid q)$.
-For a sequence of length $N$, the trajectory space grows as $|\mathcal{T}|=O(N!)$, rendering exact likelihood computation infeasible and forcing existing methods to rely on ELBO-based surrogates rather than the true objective Wang et al.
+For a sequence of length $N$, the trajectory space grows as $|\mathcal{T}|=O(N!)$, rendering exact likelihood computation infeasible and forcing existing methods to rely on ELBO-based surrogates rather than the original objective Wang et al.
 (2025a); Ou et al.
@@
 Even with an accurate likelihood approximation, a more subtle issue persists.
-In practice, rollout samples are produced by heuristic-guided policies $o\sim\pi_{\theta}^{\text{heur}}(o\mid q)$ to explore the combinatorial space.
-However, the ELBO objective still targets the likelihood of the original model distribution $\pi_{\theta}(o\mid q)$, rather than that of the heuristic-guided sampler $\pi_{\theta}^{\text{heur}}(o\mid q)$, leading to a critical mismatch between sampling and optimization that can degrade performance Schulman et al.
+In practice, rollout samples are typically produced by confidence-based generation $o\sim\pi_{\theta}^{\text{conf}}(o\mid q)$ to navigate the combinatorial space.
+However, the ELBO objective still targets the likelihood of the original model distribution $\pi_{\theta}(o\mid q)$, rather than that of the actual sampling policy $\pi_{\theta}^{\text{conf}}(o\mid q)$, leading to a mismatch between rollout and optimization that can degrade performance Schulman et al.
 (2015).
@@
 Formulation.
-Standard GRPO assumes a policy $\pi(o_{t}|o_{<t},q)$ accepting a partial sequence $o_{<t}$ with all tokens observed and predicts one token $o_{t}$ at a time, where $q$ is the query.
+Standard GRPO assumes a policy $\pi(o_{k}|o_{<k},q)$ accepting a partial sequence $o_{<k}$ with all tokens observed and predicts one token $o_{k}$ at a time, where $q$ is the query.
 Diffusion language models, however, are architected as sequence-level denoisers that accept a full sequence with mixed observed and masked tokens and predict the original values for all masked positions simultaneously.
 By forgoing arbitrary-order generation, we are able to bridge the above gap and rigorously define an AR policy $\pi_{\theta}^{\text{AR}}$ for dLLMs.
-To obtain the probability of the next token $o_{t}$ given history $o_{<t}$, we construct an input state $\tilde{x}_{t}$ where the past is observed and the future is masked:
-$\tilde{x}_{t}=[\underbrace{o_{1},\dots,o_{t-1}}_{\text{Observed}},\underbrace{\texttt{[MASK]},\dots,\texttt{[MASK]}}_{\text{Masked}}].$ (3)
-Although the diffusion language model outputs predictions for all masked positions, the autoregressive policy concerns only the next token $o_{t}$.
-We thus define $\pi_{\theta}^{\text{AR}}(o_{t}|o_{<t},q)$ as the probability distribution of $o_{t}$:
-$\pi_{\theta}^{\text{AR}}(o_{t}|o_{<t},q)\triangleq\text{Softmax}(f_{\theta}(\tilde{x}_{t}))_{t},$ (4)
-where $f_{\theta}$ denotes the model logits. Consequently, the likelihood of a complete reasoning chain $o$ is exactly computable as:
-$\pi_{\theta}^{\text{AR}}(o|q)=\prod_{t=1}^{L}\pi_{\theta}^{\text{AR}}(o_{t}|o_{<t},q).$ (5)
+To obtain the probability of the next token $o_{k}$ given history $o_{<k}$, we construct $\tilde{x}_{k}$ where the past is observed and the future is masked:
+$\tilde{x}_{k}=[\underbrace{o_{1},\dots,o_{k-1}}_{\text{Observed}},\underbrace{\texttt{[MASK]},\dots,\texttt{[MASK]}}_{\text{Masked}}].$ (5)
+Although the diffusion language model outputs predictions for all masked positions, the autoregressive policy concerns only the next token $o_{k}$.
+We define $\pi_{\theta}^{\text{AR}}(\cdot|o_{<k},q)$ as follows:
+$\pi_{\theta}^{\text{AR}}(\cdot|o_{<k},q)\triangleq\text{Softmax}(f_{\theta,k}(\tilde{x}_{k},q)),$ (6)
+where $f_{\theta,k}$ denotes the model logits at position $k$. Consequently, the likelihood of a complete reasoning chain $o$ is exactly computable as:
+$\pi_{\theta}^{\text{AR}}(o|q)=\prod_{k=1}^{|o|}\pi_{\theta}^{\text{AR}}(o_{k}|o_{<k},q).$ (7)
 Optimization.
 The above formulation enables the direct application of standard GRPO to diffusion language models.
-For each query $q$, we sample a group of outputs $\{o^{i}\}_{i=1}^{G}$ using the old policy $\pi_{\theta_{\text{old}}}^{\text{AR}}$.
+For each query $q$, we sample a group of outputs $\{o_{i}\}_{i=1}^{G}$ using the old policy $\pi_{\theta_{\text{old}}}^{\text{AR}}$.
 The objective is:
-$\mathcal{J}(\theta)$ $=\mathbb{E}_{q\sim P(Q),\{o^{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}^{\text{AR}}}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o^{i}|}\sum_{t=1}^{|o^{i}|}\left(\min\left(\rho_{i,t}\hat{A}_{i,t},\text{clip}(\rho_{i,t},1-\varepsilon,1+\varepsilon)\hat{A}_{i,t}\right)-\beta\mathbb{D}_{\text{KL}}\right)\right],$ (6)
-where $\rho_{i,t}=\frac{\pi_{\theta}^{\text{AR}}(o^{i}_{t}\mid o^{i}_{<t},q)}{\pi_{\theta_{\text{old}}}^{\text{AR}}(o^{i}_{t}\mid o^{i}_{<t},q)}$ is the probability ratio between the current and old policies and $\hat{A}_{i,t}$ denotes the group-standardized advantage.
+$\mathcal{J}(\theta)$ $=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}^{\text{AR}}}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\left(\min\left(\rho_{i,k}\hat{A}_{i,k},\text{clip}(\rho_{i,k},1-\varepsilon,1+\varepsilon)\hat{A}_{i,k}\right)-\beta\mathbb{D}_{\text{KL}}\right)\right],$ (8)
+where $\rho_{i,k}=\frac{\pi_{\theta}^{\text{AR}}(o_{i,k}\mid o_{i,<k},q)}{\pi_{\theta_{\text{old}}}^{\text{AR}}(o_{i,k}\mid o_{i,<k},q)}$ is the probability ratio between the current and old policies and $\hat{A}_{i,k}$ denotes the group-standardized advantage.
 Remarks.
@@
 The AR constraint is applied only during training to correctly assign credit.
-It refines the model’s joint distribution $p(o)$ without altering the underlying architecture.
+It refines the model’s joint distribution without altering the underlying architecture.
 At inference time, the model retains its conditional independence properties, still allowing us to employ parallel samplers Ben-Hamu et al.
 (2025) to accelerate decoding.
-JustGRPO thus achieves the reasoning depth of autoregressive models while preserving the inference speed of dLLMs (see Section 5.2).
+JustGRPO thus benefits from reasoning exploration of AR order while preserving the inference speed of dLLMs (see Section 5.2).
 5 Experiments
@@
 We apply JustGRPO on LLaDA-Instruct Nie et al.
-(2025) and evaluate its effectiveness on four standard reasoning and coding benchmarks: GSM8K, MATH, HumanEval, and MBPP.
+(2025) and evaluate its effectiveness on four standard reasoning and coding benchmarks: GSM8K, MATH-500, HumanEval, and MBPP.
 For mathematical tasks, we train on the official training split of each dataset following Zhao et al.
@@
 5.1 Main Results
-Table 1 reports the system-level comparison.
-We observe that simplifying the training objective to a standard autoregressive formulation yields consistent improvements over methods specifically designed for dLLMs.
+Table 1 reports the system-level comparison. We observe that simplifying the RL objective to a standard GRPO formulation yields consistent improvements over methods specifically designed for dLLMs.
 Performance on reasoning tasks.
@@
 These results challenge the prevailing assumption that RL for diffusion models requires optimizing over the full combinatorial space of denoising trajectories.
-Instead, treating the dLLM as a sequential generator during training appears more effective for credit assignment in logic-heavy tasks.
+Instead, treating the dLLM as a sequential generator during training appears more effective for credit assignment in general reasoning tasks.
 Robustness across generation budgets.
@@
 However, we observe robust performance gains across varying sequence lengths (128, 256, 512).
-This stability suggests that the policy has improved its fundamental reasoning capability, i.e., the ability to navigate logical branches, rather than merely memorizing fixed-length patterns.
-Table 1: System-level comparison.
-RL post-training approaches on LLaDA-Instruct.
-JustGRPO consistently achieves state-of-the-art performance across all tasks and sequence lengths.
+This stability suggests that the policy has improved its underlying reasoning capability rather than merely memorizing fixed-length patterns.
+Table 1: System-level comparison of RL post-training approaches on LLaDA-Instruct.
+JustGRPO achieves better performance than existing diffusion-specific RL methods.
 LLaDA-1.5 and LLADOU are excluded from the comparison as LLaDA-1.5 is trained on a privately collected dataset at a significantly larger scale, while LLaDOU modifies the base architecture with an auxiliary module.
@@
 Model / Seq Len 128 256 512 128 256 512 128 256 512 128 256 512
-LLaDA-1.5 - 83.3 - - - - 29.3 39.6 51.9 39.6 39.9 38.8
-LLaDOU - 88.1 - - 41.1 - - 59.1 - - 51.6 -
-D1 73.2 81.1 82.1 33.8 38.6 40.2 - - - - - -
-WD1 77.2 80.8 82.3 33.3 34.4 39.0 - - - - - -
-$d$-TreeRPO - 81.2 82.6 - 37.7 38.9 - - - - - -
-ESPO 80.0 82.3 83.7 36.0 39.0 43.4 28.1 42.1 50.0 47.4 44.6 44.2
-GDPO 78.4 82.8 84.5 33.2 39.6 41.4 26.2 39.6 39.0 43.6 50.6 47.1
-SPG 78.5 86.1 84.5 33.4 40.0 41.8 - - - - - -
-JustGRPO (Ours) 83.8 89.1 89.8 39.0 45.1 45.2 37.8 49.4 48.7 50.6 52.4 49.0
+LLaDA-1.5 Zhu et al. (2025) - 83.3 - - - - 29.3 39.6 51.9 39.6 39.9 38.8
+LLaDOU Huang et al. (2025b) - 88.1 - - 41.1 - - 59.1 - - 51.6 -
+d1 Zhao et al. (2025) 73.2 81.1 82.1 33.8 38.6 40.2 - - - - - -
+wd1 Tang et al. (2025) 77.2 80.8 82.3 33.3 34.4 39.0 - - - - - -
+$d$-TreeRPO Pan et al. (2025) - 81.2 82.6 - 37.7 38.9 - - - - - -
+ESPO Ou et al. (2025) 80.0 82.3 83.7 36.0 39.0 43.4 28.1 42.1 50.0 47.4 44.6 44.2
+GDPO Rojas et al. (2025) 78.4 82.8 84.5 33.2 39.6 41.4 26.2 39.6 39.0 43.6 50.6 47.1
+SPG Wang et al. (2025a) 78.5 86.1 84.5 33.4 40.0 41.8 - - - - - -
+JustGRPO 83.8 89.1 89.8 39.0 45.1 45.2 37.8 49.4 48.7 50.6 52.4 49.0
 5.2 JustGRPO Preserves Parallel Decoding
@@
 (2025) for parallel decoding.
-We further investigate whether the AR training constraint compromises the model’s inherent diffusion capabilities.
+Generation length is set to 256.
+We further investigate whether the AR training compromises dLLMs’ parallel decoding capabilities.
 We employ the training-free Entropy Bounded (EB) Sampler Ben-Hamu et al.
 (2025) to evaluate inference performance under varying degrees of parallelism (tokens per step).
-Figure 7 reveals that our model not only retains full compatibility with parallel decoding but exhibits a strictly superior trade-off between speed and accuracy.
-Surprisingly, the performance gain becomes more pronounced as parallelism increases.
-As shown in the MBPP and HumanEval results, while the baseline’s performance degrades sharply with more aggressive parallel steps, our model maintains stability.
+Figure 7 shows that our model retains full compatibility with parallel decoding and exhibits a strictly superior trade-off between speed and accuracy than the original LLaDA-Instruct model.
+Interestingly, the performance gain becomes more pronounced as parallelism increases.
+While the baseline’s performance degrades sharply with high parallel steps, our model maintains relatively more stable.
 For instance, on MBPP, the accuracy gap expands from +10.6% at conservative settings (1 token/step) to +25.5% at aggressive settings ($\sim$5 tokens/step).
 This observation suggests that JustGRPO does not merely fit a specific decoding path; rather, it learns a more robust reasoning manifold that is resilient to the approximation errors inherent in parallel sampling.
-The AR formulation thus acts as an effective training scaffold that refines the joint distribution $p(o)$ during optimization, providing a more stable foundation for parallel samplers to exploit at inference.
+In other words, the AR formulation acts as a training “scaffold” that refines the joint distribution $p(o)$ more effectively, providing a more stable foundation for inference.
 6 Related Work
@@
 (2025); Kim et al.
-(2025), recent research has begun to investigate its value in standard reasoning domains.
+(2025), recent research has begun to investigate its value in general reasoning domains.
 One line of research, exemplified by Dream-Coder Xie et al.
@@
 By allowing the model to bypass high-entropy tokens, effectively skipping the most demanding logical branches, arbitrary-order generation acts as an exploitation mechanism that prioritizes greedy optimization of individual trajectories at the expense of broader solution coverage.
-Therefore, eliciting the reasoning capability of dLLMs can be simpler.
+Therefore, the elicitation of reasoning capability for dLLMs can be simpler.
 By operating dLLMs in a standard autoregressive manner, we enable the direct application of Group Relative Policy Optimization (GRPO) without any complex adaptations tailored for order arbitrariness.
-This intentional constraint paradoxically yields a significant upgrade in reasoning performance, while fully preserving the parallel decoding capabilities of dLLMs.
-By returning to the basic, natural left-to-right order of language modeling, we hope to encourage a re-examination of its real value in the training of next-generation diffusion models.
+This intentional constraint paradoxically yields a significant improvement in reasoning performance, while fully preserving the parallel decoding capabilities of dLLMs.
+By returning to the basic, natural left-to-right order of language modeling, we hope to encourage a re-examination of its real value in the development of next-generation language models.
 Appendix
@@
     • $0.0$: Failure to generate a valid Markdown code block.
-Appendix B More Analysis Results
-To further validate the robustness of our findings in Section 3, we conduct extended analyses on the HumanEval benchmark using the LLaDA-Instruct model Nie et al.
+Appendix B More Analysis Results on Reasoning Boundary
+To further validate the robustness of our findings about reasoning boundary in Section 3, we conduct extended analyses on the HumanEval benchmark using the LLaDA-Instruct model Nie et al.
 (2025).
@@
 (b) Correlation between different sampling algorithms and AR in per-problem accuracy.
-We compare different sampling algorithms in pass@k performance and correlation with AR in per-problem accuracy.
+Algorithms with higher Pass@$k$ performance tend to show higher correlation with AR in problem-level performance.
 The results show that although more sophisticated sampling algorithms can achieve better Pass@$k$ performance than default confidence-based sampling, they still cannot catch up with AR order.
@@
 We observe that algorithms with higher scaling potential (higher Pass@$128$) consistently show stronger correlation with AR, with the most effective method (Neg-Entropy) achieving a correlation of $0.970$.
-This suggests that sampling algorithms with better pass@$k$ tend to behave more like AR in terms of task-level performance characteristics.
+This suggests that sampling algorithms with better Pass@$k$ tend to behave more like AR in performance characteristics.
 B.3 Block size analysis.
@@
 Since smaller block sizes explicitly restrict the model’s order flexibility, this trend is aligned with our finding that less order flexibility leads to better reasoning potential.
-B.4 Entropy Comparison Results on More Forking Tokens
+Appendix C More Results on Entropy Degradation
 Figure 11: Entropy comparison results on more forking tokens.
@@
 The specific forking tokens evaluated in this extended experiment include: “Therefore”, “Thus”, “So”, “Since”, “When”, “Given”, “However”, “Let”, “First”, “Then”, “Next”, “Finally”, “Now”, “Similarly”, “Calculate”, “Solving”, “Notice”, “Specifically”, “Follows”, “Because”, “But”, “Or”, “Consider”, “Also”, “Express”, and “Write”.
-B.5 Training Efficiency Analysis
+Appendix D Training Efficiency Analysis of JustGRPO
 […]
@@
 (2025), a representative approximation-based baseline.
-It can be observed that ESPO saturates early, effectively plateauing around 10 hours.
+It can be observed that ESPO saturates earlier, effectively plateauing around 10 hours.
 In contrast, JustGRPO exhibits a superior efficiency frontier: it surpasses the baseline’s peak accuracy in less time and maintains a continuous scaling trend.
@@
 (2025c), we demonstrate that JustGRPO allows for further acceleration via a simple heuristic: restricting gradient computation to the top-25% high-entropy tokens (Ours-Fast).
-This modification yields even faster convergence without compromising the final performance.
+This modification yields faster convergence and even slightly better performance.
 We prioritize simplicity in the current implementation and leave further efficiency optimizations for future work.

### v3

@@
 Diffusion language models.
-Inspired by the success of diffusion models in continuous image domains Ho et al.
+Motivated by the success of diffusion models in continuous domains Ho et al.
 (2020); Rombach et al.
 (2022), recent work has extended diffusion to discrete text generation.
-Early approaches operating in continuous embedding spaces Li et al.
+Early embedding-space approaches Li et al.
 (2022); Gong et al.
 (2022); Han et al.
-(2022) suffered from optimization and discretization issues.
-In contrast, masked diffusion models Lou et al.
+(2022) faced optimization and discretization challenges, leading to the adoption of masked diffusion models Lou et al.
 (2023); Sahoo et al.
@@
 (2024); Ou et al.
-(2024), which define the diffusion process directly in the discrete token space via random masking, have become the dominant paradigm.
-Notably, recent large-scale models such as LLaDA Nie et al.
+(2024), which operate directly in the discrete token space via random masking.
+Recent large-scale models like LLaDA Nie et al.
 (2025) and Dream Ye et al.
-(2025) demonstrate scalability and performance competitive with autoregressive (AR) models.
-The emergence of such large-scale diffusion language models has spurred a growing body of follow-up studies exploring their distinctive modeling and inference properties.
-Among various advantages discussed in the literature, two aspects have received particular attention.
+(2025) achieve performance competitive with autoregressive (AR) models.
+Among various advantages of dLLMs discussed in the literature, two aspects have received particular attention.
 First, dLLMs naturally support parallel decoding, enabling significant inference acceleration Wu et al.
@@
 The value of order arbitrariness.
-While early studies validated the utility of arbitrary order generation in constrained tasks like Sudoku and Zebra Puzzles Ye et al.
+While early studies validated arbitrary order generation in constrained tasks Ye et al.
 (2024); Nie et al.
 (2025); Kim et al.
-(2025), recent research has begun to investigate its value in general reasoning domains.
-One line of research, exemplified by Dream-Coder Xie et al.
-(2025) and DiffuCoder Gong et al.
-(2025), shows that diffusion models can naturally exhibit non-standard, human-like decoding behaviors (e.g., sketch-first reasoning) without any explicit supervision on generation order.
-DiffuCoder Gong et al.
-(2025) further quantifies this phenomenon by demonstrating that higher sampling temperatures reduce AR-ness, with increased order randomness correlating with improved output diversity.
-Another line of research, exemplified by P2 Peng et al.
-(2025) and LLaDOU Huang et al.
-(2025b), explicitly optimizes the decoding strategy for better generation order, thus better performance.
-Despite these advances, it remains unclear whether the observed gains primarily arise from better exploitation of existing solution patterns encoded in the data and model, or whether order arbitrariness itself enables qualitatively new reasoning strategies that are unattainable under a purely autoregressive decoding regime.
+(2025), recent research extends this to general reasoning.
+Some works demonstrate that dLLMs can naturally exhibit non-standard decoding behaviors (e.g., sketch-first) Xie et al.
+(2025); Gong et al.
+(2025), while others explicitly optimize decoding schedules to boost performance Peng et al.
+(2025); Huang et al.
+(2025b).
+Moreover, some studies Gong et al.
+(2025); Hong et al.
+(2025); Shen et al.
+(2026) also employ Pass@$k$ to demonstrate the reasoning potential of dLLMs.
+Recognizing this potential, these works predominantly focus on enhancing performance within the arbitrary-order paradigm, while leaving whether the benefit is uniquely tied to the order arbitrariness less-examined.
+Recently, Du et al.
+(2025) challenged the necessity of arbitrary order, arguing that forcing models to learn uniform permutations causes a significantly looser approximation of the underlying data distribution.
+While their analysis focuses on pre-training, we identify a parallel failure mode in reasoning and RL, revealing how the flexibility degrades exploration essential for RL training.
 Reinforcement learning for diffusion language models.
@@
 LLaDOU Huang et al.
-(2025b) explicitly models token position selection via an auxiliary policy, enabling direct estimation of trajectory likelihoods, while TraceRL Wang et al.
-(2025d) aligns optimization with inference traces through a shrinkage-based step-wise MDP formulation.
-Nevertheless, these approaches remain committed to preserving the full arbitrary-order generation mechanism, implicitly treating its structural complexity as indispensable, whereas we question whether effective RL training can be achieved more simply by reexamining the necessity of arbitrary order itself.
+(2025b) introduces an auxiliary policy to model token position selection, enabling direct trajectory likelihood estimation, while TraceRL Wang et al.
+(2025d) aligns optimization with inference traces via a shrinkage-based step-wise MDP.
+Nevertheless, both retain the full arbitrary-order mechanism, implicitly assuming its necessity, whereas we question whether effective RL training can be achieved more simply by reconsidering the need for arbitrary order itself.
 7 Conclusion